What are the three consecutive odd integers whose sum is 521?
Answer by stanbon(48570)

What are the three consecutive odd integers whose sum is 521?
------------------------
1st 2x-1
2nd 2x+1
3rd 2x+3
------------------
Equation:
6x+3 = 521
6x = 518
x = 56 1/3
-------------------
1st: 2x-1 = 3(56 1/3)-1 is not an integer
================================================
Cheers,
Stan H.